OpenStudy (anonymous):

What is -1 plus x divided by 2 and equals 3?

OpenStudy (anonymous):

Do you mean this equation? \[\frac{ (-1+x) }{ 2 }=3\]

OpenStudy (anonymous):

Yes that equation

OpenStudy (anonymous):

 \[x-1=6\]\[x=7\]

OpenStudy (anonymous):

Okay thank you. Can you answer one more equation for me?